jaelaj3:

Which equation does the graph below represent? A coordinate grid is shown. The x axis values are from negative 5 to positive 5 in increments of 1 for each grid line, and the y axis values are from negative 10 to positive 10 in increments of 2 for each grid line. A line is shown passing through the ordered pairs negative 4, 8 and 0, 0 and 4, negative 8.

snowflake0531:

so, you know that the points are (4,8), (0,0), and (4,-8) but then that wouldn't be a linear equation... could you please screenshot the question?

jaelaj3:

1 attachment
snowflake0531:

so can i just ignore the other stuff that you said, and just look at the graph and find the equation?

jaelaj3:

y = −2x y = fraction negative 1 over 2x y = 2x y = fraction 1 over 2x those are the answeres

jaelaj3:

yes

snowflake0531:

okay, do you know how to find the slope of an equation? or the y-intercept?

jaelaj3:

rise over run

snowflake0531:

yes so, choose 2 random coordinate points for me and write them

jaelaj3:

1.-2

snowflake0531:

okay, and anotehr one?

snowflake0531:

and it would be easier if they were closer together by the way

jaelaj3:

2. -4

snowflake0531:

okay, and by teh way, coordinates are written as (1,-2) and (2,-4) okay, so do you know the actual formula to find the slope, not just rise over run?

jaelaj3:

y= +x

jaelaj3:

or times x

jaelaj3:

this is the formula (y2-y1)/(x2-x1)

snowflake0531:

yes, do you know hwo to use it?

jaelaj3:

take the points and use the eqaution (y2-y1) (4-0)= 4 (x2-x1) (8-0)= 8 so the slope is equal to 4/8 4/8 simplifies to 1/2 so the slope is 1/2x

jaelaj3:

i just need to know if its positive or negative

jaelaj3:

and why

snowflake0531:

ummm, that's not really how you do it

jaelaj3:

ok, where am I wrong?

jaelaj3:

idk

snowflake0531:

okay, so, y2-y1/x2-x1 in (1,-2) and (2,-4) 2 and 1 can be our x and y 1's, and -4 and -2 can be our x and y 2nd's so, give me a sec, i'll draw it out

snowflake0531:

 if that makes sense

jaelaj3:

the answer?

jaelaj3:

little confused

jaelaj3:

so its y =1 over 2x

snowflake0531:

so -4 - (-2) = -2 2-1 =1 -2/1 =-2

snowflake0531:

*-2x